OpenStudy (anonymous):

One milliliter of blood contains approximately 5 million red blood cells, and one pint is about 473 milliliters. Create and solve an equation that can be used to find approximately how many red blood cells are in a pint of blood. Use scientific notation for the first box of the equation. + 5 × 109 2.4 × 109 473 × 5 × 106 9.73 × 109 4.73 – ÷ 9.73 × 106 2.4 × 106

OpenStudy (anonymous):

Hi catwomen12 --- don't know if you've got this one done yet. You might consider unit analysis: there's 1 ml blood per 5x10^6 red blood cells; view this as the fraction \[ \frac{1\text{ml blood}}{5\times 10^6 \text{ml red blood cells}}\] Then, since 1 pt = 473 ml, the fraction \[\frac{1 \text{pt}}{473 \text{ ml blood}}\] is just the number 1!! (Numerator and denominator are the same.) If you multiply these two fractions together, you're really just *re-naming" the first fraction (since you're multiplying it by one); what cancels out? What's left?